# LinPrim: Linear Primitives for Differentiable Volumetric Rendering

### v1

@@ 4. Experiments
While there is a large body of work offering improvements to 3DGS, we mainly focus on the original work and the changes made by Mip-Splatting, as they most closely relate to the foundational changes made by our approach. We believe that much of the recent progress made on improving 3DGS can similarly be applied to our approach to further improve quality or performance metrics.
4.1. Reconstruction Quality
39f36da05b 5a269ba6fe dc263dfbf0 08bbbdcc3d fb564c935d PSNR SSIM LPIPS Primitives
[…]
Table 1. Quantitative Results on ScanNet++ v2 scenes (Yeshwanth and Liu et al., 2023). We limit GS-MCMC (Kheradmand et al., 2024) to use as many primitives as our approach does on average. Our method generates results of comparable performance while using much fewer primitives than 3DGS (Kerbl et al., 2023) and Mip-Splatting (Yu et al., 2024).
@@ 4.1. Reconstruction Quality
Zip-NeRF (2023) 28.54 0.828 0.189
Instant-NGP (2022) 25.68 0.705 0.302
[…]
Table 2. Quantitative Results on Mip-NeRF 360 scenes (Barron et al., 2022). Our approach generates results comparable in performance while only using $54\%$ of the primitives of 3DGS and $43\%$ of Mip-Splatting.
We measure reconstruction quality using the standard image metrics PSNR, SSIM and LPIPS. On the ScanNet++ dataset, our method consistently achieves scores close to both 3DGS and Mip-Splatting despite using significantly fewer primitives, as visible in Table 1. We also compare against GS-MCMC (Kheradmand et al., 2024), which allows precise limitations to the primitive population, and find that our results remain comparable, albeit slightly worse. For Mip-NeRF 360 scenes, we see a similar trend of comparable performance but with a more compact representation. Figures 6 and 7 illustrate that our method faithfully reconstructs many challenging regions while sometimes showing slightly lower overall fidelity. These outcomes stem, in part, from the bounded nature of our primitives, which enforce sharper, more “binary” decisions in underconstrained areas. In practice, this can lead to crisper reconstructions in reflective or transparent regions—such as glass or distant backgrounds—but can also result in more conspicuous edges on rarely observed surfaces like ceilings. Overall, we find our approach to work the best in smaller, more densely captured scenes, as it retains strong geometric and visual clarity without overly densifying frequently observed regions.
4.2. Tetrahedra Evaluation
In Table 3, we evaluate our tetrahedron approach using the same datasets, parameter constraints, and optimization framework described previously. By simply swapping octahedra for tetrahedra, we demonstrate that our method generalizes well to polyhedra with different geometric constraints and can still produce photorealistic reconstructions, as seen in Figure 5. Quantitatively, our tetrahedron approach performs similarly to octahedra on the ScanNet++ scenes and struggles on Mip-NeRF 360, highlighting the benefits of the increased symmetry and stability inherent to our octahedra.
Figure 5. Qualitative Results of our Tetrahedron approach on the Mip-NeRF 360 (2022) Bicycle scene and the ScanNet++ v2 (2023) 39f36da05b scene.
@@ 4.2. Tetrahedra Evaluation
[…]
Table 3. Comparison of the performance of our Octahedron and Tetrahedron approaches on ScanNet++ v2 (2023) and Mip-NeRF 360 (2022) scenes. Both approaches produce comparable results on ScanNet++ scenes, but our Octahedron approach reaches higher reconstruction quality on Mip-NeRF 360 scenes.
Although our existing pipeline was designed primarily around octahedra, these results confirm that it is not narrowly tailored to one specific primitive type. Nonetheless, tetrahedra could likely benefit from specialized adaptive population control and projection techniques that fully consider their asymmetric geometry, potentially closing the performance gap on currently challenging scenes. For instance, tuning how we split tetrahedra could yield further gains in accuracy.
4.3. Rendering Speed
@@
[…]
Table 4. Quantitative Results on the rendering speed of tested approaches. We report the average time to render every train and test view on ScanNet++ v2 and Mip-NeRF 360 scenes.
@@
These results confirm that the majority of the rendering cost stems from iterating over each face of each primitive to compute ray intersections. Since tetrahedra contain fewer faces, each pass requires less computational effort, thereby boosting the frame rate substantially. While our methods consistently demonstrate real-time rendering across both datasets, we must consider that the difference in primitive count can impact performance. Nevertheless, these experiments underscore the capability of our linear primitives, whether octahedra or tetrahedra, to balance high-fidelity reconstruction with the interactivity needed for real-time applications.
4.4. Anti-Aliasing Ablation
@@
[…]
Table 5. Ablation showing the impact of the two filters used in our anti-aliasing efforts. Combining both filters reduces the reconstruction error.
4.5. Limitations
While our method demonstrates comparable results and is easily understood and modified, it is not without drawbacks. First, the bounded nature of our primitives can introduce hard edges in poorly observed regions, resulting in more ”segment-like” artifacts under limited view coverage. In these regions, the ability of Gaussian kernels to smoothly transition into each other leads to more visually pleasing results, even if the reconstruction is not necessarily better.
Second, the current population control processes are not as consistent as their Gaussian counterparts. While we did adapt the processes to our primitives, the underlying paradigms are not as suitable for our polyhedron primitives. More specifically, polyhedra could benefit from more specialized processes for placing newly split primitives and improved criteria for determining which areas should be densified. It would be especially interesting to explore enhancements to 3DGS that address the above issues, as such improvements could disproportionately benefit our polyhedron-based approaches.
5. Conclusion
Promising future work could focus on leveraging the inherent properties of our representation. As a concrete example, while the rendering improvements made by EVER (Mai et al., 2024) create a lot of computational overhead for Gaussian-based approaches, the core requirements - discrete bounds and homogeneous density - are already fulfilled by our primitives. A different avenue would be the incorporation of improved population control mechanisms. To summarize, we hope our linear primitives open up new possibilities for representing scenes in NVS applications and foster the creation of novel scene representations. Ultimately, the versatility of our approach presents a promising opportunity for representations that bridge the gap between NVS and the real world.

### v3

@@ 4 Experiments
While there is a large body of work offering improvements to 3DGS, we mainly focus on the original work and the changes made by Mip-Splatting, as they most closely relate to the foundational changes made by our approach. We believe that much of the recent progress made on improving 3DGS can similarly be applied to our approach to further improve quality and performance metrics.
Figure 4: Qualitative Results on test views from ScanNet++ v2 scenes [29].
4.1 Reconstruction Quality
We measure reconstruction quality using the standard image metrics PSNR, SSIM and LPIPS. On the ScanNet++ dataset, our method consistently achieves scores close to both 3DGS and Mip-Splatting despite using significantly fewer primitives, as visible in Table 1. We also compare against GS-MCMC [14], which allows precise limitations to the primitive population, and find that our results remain comparable, albeit slightly worse. For Mip-NeRF 360 scenes, we see a similar trend of comparable performance but with a more compact representation, as can be seen in Table 2. These results align with other works that suggest that the number of used Gaussians can be drastically reduced without a negative impact on performance through introduction of more advanced population control paradigms [8, 9]. While our approach introduces fundamental changes, we deliberately preserve components not directly related to the primitives. As such, more advanced paradigms can be applied in our primitives in a similar manner.
Figure 4 illustrates that our method faithfully reconstructs many challenging regions while sometimes showing slightly lower overall fidelity. These outcomes stem, in part, from the bounded nature of our primitives, which enforce sharper, more “binary” decisions in underconstrained areas. In practice, this can lead to crisper reconstructions in reflective or transparent regions—such as glass or distant backgrounds—but can also result in more conspicuous edges on rarely observed surfaces like ceilings. Overall, we find our approach to work the best in smaller, more densely captured scenes, as it retains strong geometric and visual clarity without overly densifying frequently observed regions. Further results on Mip-NeRF 360 scenes can be seen in Appendix C.
Table 1: Quantitative Results on ScanNet++ v2 scenes [29]. We limit GS-MCMC to use as many primitives as our approach does on average. Our method generates results of comparable performance while using much fewer primitives than 3DGS and Mip-Splatting.
PSNR 39f36da05b 5a269ba6fe dc263dfbf0 08bbbdcc3d fb564c935d Prim. 3DGS [13] 27.693 23.675 23.436 22.444 23.210 738k Mip-Splatting [30] 27.728 23.636 23.454 22.570 23.221 977k GS-MCMC [14] 27.902 23.787 23.782 23.657 23.371 255k LinPrim (ours) 27.557 23.845 22.668 22.312 23.801 255k
Table 2: Quantitative Results on Mip-NeRF 360 scenes [2]. Our approach generates results comparable in performance while using fewer primitives.
@@ 4.1 Reconstruction Quality
Zip-NeRF [3] 28.54 0.828 0.189
Instant-NGP [20] 25.68 0.705 0.302
3DGS [13] 27.43 0.813 0.217 3.32M
Mip-Splatting [30] 27.79 0.827 0.203 4.17M
LinPrim (ours) 26.63 0.803 0.221 1.79M
4.2 Tetrahedra Evaluation
In Table 3, we evaluate our tetrahedron approach using the same datasets, parameter constraints, and optimization framework described previously. By simply swapping octahedra for tetrahedra, we demonstrate that our method generalizes well to polyhedra with different geometric constraints and can still produce photorealistic reconstructions. Quantitatively, our tetrahedron approach performs similarly to octahedra on the ScanNet++ scenes and struggles on Mip-NeRF 360, highlighting the benefits of the increased symmetry and stability inherent to our octahedra. Although our existing pipeline was designed primarily around octahedra, these results confirm that it is not narrowly tailored to one specific primitive type. Nonetheless, tetrahedra could likely benefit from specialized adaptive population control and projection techniques that fully consider their asymmetric geometry, potentially closing the performance gap on currently challenging scenes. For instance, tuning how we split tetrahedra could yield further gains in accuracy.
Table 3: Comparison of the performance of our Octahedron and Tetrahedron approaches on ScanNet++ v2 [29] and Mip-NeRF 360 [2] scenes. Both approaches produce comparable results on ScanNet++ scenes, but our Octahedron approach reaches higher reconstruction quality on Mip-NeRF 360 scenes.
@@ 4.2 Tetrahedra Evaluation
Octahedron 24.04 0.849 0.281 26.63 0.803 0.221
Tetrahedron 24.05 0.848 0.302 25.96 0.790 0.247
4.3 Limitations
While our method demonstrates comparable results and is easily understood and modified, it is not without drawbacks. First, the bounded nature of our primitives can introduce hard edges in poorly observed regions, resulting in more "segment-like" artifacts under limited view coverage. In these regions, the ability of Gaussian kernels to smoothly transition into each other leads to more visually pleasing results, even if the reconstruction is not necessarily better. Second, the current population control processes are not as consistent as their Gaussian counterparts. While we did adapt the processes to our primitives, the underlying paradigms are not as suitable for our polyhedron primitives. More specifically, polyhedra could benefit from more specialized processes for placing newly split primitives and improved criteria for determining which areas should be densified. It would be especially interesting to explore enhancements to 3DGS that address the above issues, as such improvements could disproportionately benefit our polyhedron-based approaches.
Broader Impact.
Our work provides a new approach to novel view synthesis, which extends the design space of 3D scene representations. It has the potential to improve the quality and performance of 3D reconstruction and downstream applications, benefiting research in graphics, vision, robotics, and related fields. We do not anticipate any adverse environmental, societal or ethical effects.
5 Conclusion
Promising future work could focus on leveraging the inherent properties of our primitives. As a concrete example, their triangular faces closely resemeble traditional meshes. Developing efficient and accurate methods to binarize primitive opacity [7] presents significant potential for directly optimizing mesh representations from images. Additionally, although the rendering enhancements introduced by EVER [17] require significant changes for Gaussian-based methods, our primitives inherently allow for easily finding intersections with homogeneous primitives, avoiding the need for ray-tracing and enabling exact blending without additional computational overhead. Finally, integrating more specialized or advanced population control mechanisms could significantly enhance performance by improving primitive density in optimized scenes.
To summarize, we hope our linear primitives open up new possibilities for representing scenes in NVS applications and foster the creation of novel scene representations.
Acknowledgements.
This work was supported by the ERC Consolidator Grant Gen3D (101171131) and the German Research Foundation (DFG) Research Unit “Learning and Simulation in Visual Computing”. We thank Angela Dai for the video voice-over.
Appendix A Rendering Speed
@@
Table 4: Quantitative Results on the rendering speed of tested approaches. We report the average time to render every train and test view on ScanNet++ v2 and Mip-NeRF 360 scenes.
@@
Tetrahedron (ours) 5.7ms 15.2ms
@@
These results confirm that the majority of the rendering cost stems from iterating over each face of each primitive to compute ray intersections. Since tetrahedra contain fewer faces, each pass requires less computational effort, thereby boosting the frame rate substantially.
While our methods consistently demonstrate real-time rendering across both datasets, we must consider that the difference in primitive count can impact performance. Nevertheless, these experiments underscore the capability of our linear primitives, whether octahedra or tetrahedra, to balance high-fidelity reconstruction with the interactivity needed for real-time applications.
Appendix B Anti-Aliasing
@@
Table 5: Ablation showing the impact of the two filters used in our anti-aliasing efforts. Combining both filters reduces the reconstruction error.
@@
30.84 0.929 0.196 22.46 0.870 0.267
30.96 0.931 0.196 22.59 0.870 0.266
30.97 0.930 0.195 22.66 0.871 0.266
31.21 0.936 0.195 22.67 0.871 0.266
Appendix C Additional Results
In Figure 5 we show qualitative results of our representation on Mip-NeRF 360 scenes. In Table 6 we show the average PSNR, SSIM, and LPIPS metrics on the used ScanNet++ v2 scenes.
Figure 5: Qualitative Results on test views from Mip-NeRF 360 scenes [2].
Table 6: Quantitative Results on ScanNet++ v2 scenes [29]. We limit GS-MCMC to use as many primitives as our approach does on average. Presented results are the average over the five scenes considered during our evaluation.
PSNR PSNR SSIM LPIPS Primitives 3DGS [13] 24.092 0.853 0.263 738k Mip-Splatting [30] 24.122 0.852 0.261 977k GS-MCMC [14] 24.500 0.862 0.256 255k LinPrim (ours) 24.037 0.849 0.281 255k